OpenStudy (anonymous):

Between which 2 consecutive integers does the square root of 55 lie?

OpenStudy (anonymous):

can you think of a squared number thats a little higher than 55 and a little lower than 55?

OpenStudy (anonymous):

What do you mean like 50 and 60?

OpenStudy (anonymous):

list of squared integers- you should probably memorize some of these 1^2=1 2^2=4 3^2=9 4^2=16 5^2=25 6^2=36 7^2=49 8^2=64 9^2=81 10^2=100 11^2=121 12^2=144 13^2=169- i think 15^2=225 25^2=625

OpenStudy (anonymous):

Im really confused about what your saying I'm just wondering what two consecutive integers the square root of 55 lies between

OpenStudy (anonymous):

you could always use a calculator, calculate the square root of 55, and then look at the 2 numbers it lies between

OpenStudy (anonymous):

\[\sqrt{100}=10\]\[\sqrt{81}=9\]\[\sqrt{64}=8\]\[\sqrt{49}=7\]\[\sqrt{36}=6\]\[\sqrt{25}=5\]

OpenStudy (anonymous):

the square root of 55 would lie between 49 and 64 right?

OpenStudy (anonymous):

so the 2 integers it would lie between would be 7 and 8

OpenStudy (anonymous):

Thanks anyways but I already figured it out